In July of 1999 a planet was reported to be orbiting the Sun-like star Iota Horologii with a period of 320 days. Find the radius of the planet’as orbit, assuming that Iota Horologii has the same mass as the Sun. (This planet is presumably similar to Jupiter, but it may have large, rocky moons that enjoy a relatively pleasant climate.)
Answer #1

SOLUTION :

Let the radius of orbit is r

Mass of star , M = 1.99 *10^30 kg

Time period, T = (320 * 24 * 60 * 60 ) sec

As T^2/r^3 = 4 * pi^2/(G *M)

(320 * 24 * 60 * 60 )^2/(r^3) = 4*pi^2/(6.673 *10^-11 * 1.99 *10^30)

Solving for r

r = 1.37 *10^11 m

The radius of the orbit is 1.37 *10^11 m